fvbbgfvcbgvgbnvbhgfv:

Gene paid a deposit on a leased car. The deposit earns 2.8 percent simple annual interest. At the end of the year, the interest that is earned is $22.40. What was the amount of the original deposit?

Mercury:

simple interest: i = p * r * t where p = the principal (what you'll solve for), r is the interest rate as a decimal (0.028) and t = time (1 year), and i = the interest earned ($22.40) plug in and solve for p